OpenStudy (anonymous):

what is the inverse of y=x^2

OpenStudy (anonymous):

x=y^2, replace x with y, y with x

OpenStudy (anonymous):

\[y=x ^{2}\] \[x=y ^{2}\] \[\sqrt{x}=y\] Final answer is: \[y=\sqrt{x}\]

OpenStudy (anonymous):

plus or minus sqrt(x) if you were to solve for y

OpenStudy (anonymous):

I always forget that /facepalm good one rickjbr.

hero (hero):

Luis, you should have included y = f(x) first before replacing y with f(x)